How much charge is on each plate of a 4.00-µF capacitor when it is connected


to a 12.0-V battery (? (b) If this same capacitor is connected to a 1.50-V


battery, what charge is stored?


1
Expert's answer
2022-03-15T10:33:45-0400

(a) the charge on the plates of capacitor

"q=CV=4.00*10^{-6}*12.0=4.8*10^{-5}\:\rm C"

(b)

"q=CV=4.00*10^{-6}*1.50=6.0*10^{-6}\:\rm C"